What is the maximum mass of S8 that can be produced by combining 76.0 g of each reactant?

8SO2+16H2S⟶3S8+16H2O
1
Expert's answer
2020-10-05T14:23:50-0400

At first, molar masses of compounds should be found:

M(SO2)=32+2×16=64(g/mol),

M(H2S)=2×1+32=34(g/mol),

M(S8)=8×32=256(g/mol).

1)If there are 76g of SO2 mass of S8 may be found from the proportion:

x=76×3×256/8×64=114(g).

2)If there are 76g of H2S:

x=76×3×256/16×34=107.3(g).